### v1

# Analytic approach to boundary integrability with application to mixed-flux $AdS_{3}\times S^{3}$

Abstract
Boundary integrability provides rare analytic control over field theories in the presence of an interface, from quantum impurity problems to open string dynamics. We develop an analytic framework for integrable boundaries in two-dimensional sigma-models that determines admissible reflection maps directly from the meromorphic Lax connection. Applying it to open strings on $AdS_{3}\times S^{3}$ with mixed NSNS and RR flux, we find two branches of integrable boundary conditions. One branch admits D-branes wrapping twisted conjugacy classes on $SU(1,1)\times SU(2)$, with the mixed-flux deformation encoded entirely into dynamical boundary data. At the exactly solvable WZW point these coincide with the conformal D-branes, providing a natural link to conformal perturbation theory.
preprint: HU-EP-26/02-RTG
Field theories with boundaries play an important role in both high-energy and condensed-matter physics. In open string theory, worldsheet boundary conditions describe D-branes, while in condensed-matter and lattice systems they model defects and boundary critical phenomena. When the bulk theory is integrable, the situation becomes especially interesting, as an infinite set of conserved charges provides rare analytic control of the interacting quantum dynamics. The presence of a boundary probes this structure in a nontrivial way. Geometric symmetries are typically broken, and it is far from obvious that an infinite tower of conserved charges can survive. When they do, notions of boundary integrability have opened access to exact reflection data (see e.g. [1]) and methods for defect observables in AdS/CFT (see e.g. [2]). Identifying which physical interfaces are compatible with integrable dynamics is therefore a basic structural question. In this Letter, we address this for two-dimensional sigma-models from a novel perspective based on very simple analytical arguments.
The starting point is the one-parameter family of flat Lax connections ${\cal L}(\mathsf{x})$ representing the equations of motion of a periodic sigma-model on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$, i.e. [^1] For brevity, and where possible, we will suppress explicit time ($\tau\in\mathbb{R}_{\tau}:=\mathbb{R}$) or space ($\sigma\in S^{1}_{\sigma}:=S^{1}$ with radius 1) dependence. $\mathrm{d}{\cal L}(\mathsf{x})+{\cal L}(\mathsf{x})\wedge{\cal L}(\mathsf{x})=0$ and ${\cal L}(\sigma+2\pi;\mathsf{x})={\cal L}(\sigma;\mathsf{x})$ for every value of the spectral parameter $\mathsf{x}$. This ensures that the monodromy matrix $T_{\cal L}(2\pi,0;\mathsf{x})$, with
$T_{\cal L}(b,a;\mathsf{x})=\overset{\leftarrow}{\mathcal{P}}\exp\left(-\int^{b}_{a}d\sigma{\cal L}_{\sigma}(\mathsf{x})\right),~~~~a,b\in\mathbb{R},$ (1)
is a generator of conserved charges, as its eigenvalues are independent of $\tau$ for all $\mathsf{x}$ (see e.g. the lecture notes [3]).
On an interval $\sigma\in[0,\pi]$, boundary integrability requires a modification of this generator. The prototype is obtained by following Sklyanin’s construction [4, 5], where one forms a double-row monodromy matrix $T_{\text{\tiny b}}(\mathsf{x})$ by gluing $T_{\cal L}(\pi,0;\mathsf{x})$ over the interval $\sigma\in[0,\pi]$ to a reflected copy. On the lattice, Sklyanin’s approach applies to integrable models with crossing-symmetric $R$-matrices of difference form [5]. The reflection compatible with integrability then has two components: a spatial flip reversing the order of sites, and a spectral reflection $\mathsf{x}\mapsto\rho(\mathsf{x})=-\mathsf{x}$ implementing crossing symmetry. Additional boundary degrees of freedom are encoded in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ at the endpoints $\sigma_{\star}=0,\pi$, which must satisfy the Boundary Yang-Baxter Equation to preserve integrability. Translating to the sigma-model, the analogue is to apply a worldsheet parity flip $\sigma\rightarrow 2\pi-\sigma$ combined with a spectral reflection $\rho(\mathsf{x})$ that leaves the Lax connection invariant (see e.g. [4, 6, 7, 8]). One can then effectively set $T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T_{\cal L}^{-1}(\pi,0;\rho(\mathsf{x}))K_{\pi}^{-1}(\mathsf{x})T_{\cal L}(\pi,0;\mathsf{x})$ (up to automorphisms implementing other conventions) and solve the conditions for $K_{\sigma_{\star}}$ under which infinite charges are generated by $T_{\text{\tiny b}}(\mathsf{x})$.
However, there are limitations to taking this prescription directly as a definition of boundary integrability for sigma-models. On the lattice, $\rho(\mathsf{x})$ is dictated by the crossing structure of a chosen $R$-matrix, and the familiar $\rho(\mathsf{x})=-\mathsf{x}$ is tied to those of difference form. For an integrable sigma-model, such quantum input is often unavailable (or only indirect through the sigma-model $S$-matrix), and the defining object is just a Lax connection ${\cal L}(\mathsf{x})$ that generically is meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$. Crucially, there need not exist a $\rho(\mathsf{x})$ that leaves ${\cal L}(\mathsf{x})$ invariant after the parity flip, and in this case there is generally not an immediate principle that selects $\rho(\mathsf{x})$ without additional structure such as e.g. quantum crossing.
In this Letter, we instead determine $\rho(\mathsf{x})$ intrinsically from the analytic structure of the Lax connection, bypassing input on bulk parity or crossing invariance. [^2] Alternative Lax-based approaches via parity exist [9, *Corrigan:1996nt], but they are based on auxiliary constructions and order-by-order expansions in $\mathsf{x}$. Here we have a purely analytic approach, which moreover admits a natural origin in 4-dimensional Chern-Simons theory that is compatible with the sigma-model variational principle, as will be reported elsewhere [11]. The key point is that ${\cal L}(\mathsf{x})$ is defined only up to gauge transformations, so the notion of a “reflected copy” should be formulated modulo this freedom. Accordingly we allow the reflected connection to be chosen within the gauge orbit of ${\cal L}(\rho(\mathsf{x}))$, and fix both the relevant representative $U$ and $\rho(\mathsf{x})$ by demanding that the divisor of zeros and poles on $\mathbb{CP}^{1}$ is preserved under the reflection. In the double-row picture the boundary values of this bulk gauge may be packaged as reflection matrices $K_{\sigma_{\star}}$, but the conceptual shift is that we use $U$ to determine $\rho(\mathsf{x})$ from bulk analyticity prior to solving for reflection data.
As an important case, we apply this approach for string sigma-models on $AdS_{3}\times S^{3}\times M_{4}$ with mixed NSNS and RR fluxes; for reviews see [12, *Seibold:2024qkh]. The classical Lax connection of this model is explicitly known for all flux values [14] and illustrates the above limitation: the parity-based criterion alone does not identify $\rho(\mathsf{x})$ at nontrivial NSNS flux. Moreover, while worldsheet methods have led to $S$-matrices [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda], a dual CFT and lattice formulation remains absent for generic flux ratios. This makes the mixed-flux background an ideal arena to demonstrate our practical approach to boundary integrability.
Beyond such technical advantages, boundary integrability of the mixed-flux $AdS_{3}\times S^{3}$ background is also motivated by exact worldsheet studies. At the pure NSNS point, the bosonic theory is described by an exactly solvable WZW model on $G=SU(1,1)\times SU(2)$ [20, *Maldacena:2000hw], in which CFT-preserving D-branes wrap twisted conjugacy classes that admit a boundary CFT (BCFT) description [22, *Stanciu:1999id, *Figueroa-OFarrill:2000gfl, 25]. Turning on RR-flux corresponds to a marginal deformation and, for pointlike branes, its effect on the open-string spectrum has been analysed by boundary conformal perturbation theory, while in the closed-string sector the RR perturbation is singular [26]. The boundary sector thus provides one of the few controlled windows on mixed-flux physics. It was in fact suggested in [14] that identifying integrable boundary conditions could provide the missing input at finite flux from integrability and offer a bridge to the BCFT picture. The present Letter realises this at the classical bosonic level. Imposing our analytic criterion on the mixed-flux Lax yields two branches of integrable boundary conditions: one compatible with the sigma-model action only in the pure RR limit, while the other persists for generic flux and includes all D-branes wrapping twisted conjugacy classes. Interestingly, at the level of the reflected copy, the former corresponds to identifying $\rho(\mathsf{x})$ from the invariance of ${\cal L}(\mathsf{x})$ under parity alone, whereas the latter requires combining worldsheet parity with a transformation within the Lax gauge orbit. It is in the latter branch that we have a rigid class of integrable branes whose embedding is fixed by the undeformed metric. We show that their gauge-invariant worldvolume two-form can remain flux-independent, while the mixed-flux contribution is encoded entirely in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ that are dynamical. [^3] This is consistent with the $\lambda$-deformed WZW case, where integrable D-branes also remain twisted conjugacy classes but for which the gauge-invariant worldvolume two-form acquires a non-trivial $\lambda$–dependence given the $\lambda$–deformed metric [27]. This provides a natural setup for comparing integrability-based reflection data with conformal perturbation theory around the solvable WZW point, where the same branes underlie the BCFT description. In particular, it would be very interesting to include the fermionic sector, study the open spectrum and connect to [26], which we leave for future work.
I Analytic criteria
To analyse integrable two-dimensional sigma-models with boundaries, we start from bulk data: a flat Lax connection ${\cal L}(\mathsf{x})$, generally meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$ and in an $n\times n$ matrix representation of the loop algebra ${\mathbb{C}(\mathsf{x})}\otimes\mathfrak{g}$, whose analytic structure in $\mathsf{x}$ encodes an infinite tower of conserved charges. Under periodic conditions on $S^{1}_{\sigma}$, these are generated by $T_{{\cal L}}(2\pi,0;\mathsf{x})$. One can also allow for flatness-preserving gauge transformations
${\cal L}^{U}(\mathsf{x})=U{\cal L}(\mathsf{x})U^{-1}-\mathrm{d}UU^{-1},$ (2)
with any dynamical $n\times n$ matrix $U:=U(\tau,\sigma)$. Then $T_{{\cal L}^{U}}(2\pi,0;\mathsf{x})=U(\tau,2\pi)T_{\cal L}(2\pi,0;\mathsf{x})U(\tau,0)^{-1}$ also generates conserved charges when $U$ is periodic in $\sigma$.
With open boundaries, we propose the following ansatz of the double-row monodromy matrix
$T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T^{-1}_{{\cal L}^{U}}(\pi,0;\rho(\mathsf{x}))K_{\pi}(\mathsf{x})^{-1}T_{\cal L}(\pi,0;\mathsf{x}),$ (3)
which generates an infinite tower of conserved quantities under the integrable boundary conditions [^4] Flatness of the Lax and the boundary conditions (4) indeed imply that $\partial_{\tau}\mathrm{Tr}T_{\text{\tiny b}}(\mathsf{x})^{p}=0$ for any integer $p$, so that expanding the eigenvalues of $T_{\text{\tiny b}}(\mathsf{x})$ in $\mathsf{x}$ produces conserved objects.
${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}K_{\sigma_{\star}}(\mathsf{x}){\cal L}_{\tau}^{U}(\rho(\mathsf{x}))K_{\sigma_{\star}}^{-1}(\mathsf{x})-\partial_{\tau}K_{\sigma_{\star}}(\mathsf{x})K_{\sigma_{\star}}^{-1}(\mathsf{x}),$ (4)
with $\overset{\text{\tiny b}}{=}$” denoting evaluation at the boundary $\sigma_{\star}=0,\pi$ and $K_{\sigma_{\star}}(\mathsf{x}):=K(\tau,\sigma_{\star};\mathsf{x})$ is the reflection matrix.
Typically, $K_{\sigma_{\star}}(\mathsf{x})$ is an $n\times n$ matrix in $\mathbb{C}(\mathsf{x})\otimes G$, with $G$ the Lie group of $\mathfrak{g}$. One may also allow constant matrices $K=w\not\in G$ that implement (possibly outer) automorphisms of $\mathfrak{g}$ as $W(y)=wyw^{-1}\in\mathfrak{g}$ for $y\in\mathfrak{g}$ [7]. At the lattice level, such choices can correspond to different representation assignments in the reflected monodromy. In suitable matrix realisations, they can also cover the common alternative convention where the inverse in (3) is replaced by transposition.
Note that in (4) only the endpoint values $U(\tau,\sigma_{\star})$ enter, and with (2) one may thus see the bulk gauge $U$ as part of a reflection matrix $\tilde{K}=KU$. However, as discussed earlier, the role of $U$ is fundamentally distinct: it is used to select a $\rho(\mathsf{x})$ for the bulk, while we reserve the notation $K$ for genuine, $\sigma$-independent, reflection matrices. This viewpoint will let us systemise the prescription of [4, 6, 7, 8, 28, 27, 29] through model-independent arguments that are not based on parity-invariant gluing.
Fixing $U$ and $\rho(\mathsf{x})$ from analyticity.
Our main point is that $\rho(\mathsf{x})$ can be fixed intrinsically from comparing the analytic structure of ${\cal L}(\mathsf{x})$ and the reflected representative ${\cal L}^{U}(\rho(\mathsf{x}))$, mirroring the lattice logic where $\rho(\mathsf{x})$ is derived from bulk properties of the $R$-matrix (rather than boundary $K$-matrices).
A first crucial ingredient is a choice of “physical” representatives for the unreflected and reflected rows, such that $T_{\text{\tiny b}}(\mathsf{x})$ can generate a set of local charges. Concretely, we require that there is a nonsingular $\mathsf{x}_{0}\in\mathbb{CP}^{1}$ where
${\cal L}_{\sigma}(\mathsf{x}_{0})=0,\qquad{\cal L}_{\sigma}^{U}(\rho(\mathsf{x}_{0}))=0,$ (5)
so that the double-row monodromy admits a local expansion around $\mathsf{x}_{0}$. Note that $K_{\sigma_{\star}}(\mathsf{x})$ is $\sigma$-independent and thus, when regular at $\mathsf{x}_{0}$, does not alter locality.
Let us now denote by $\mathrm{D}_{\text{\tiny p}}[{\cal L}(\mathsf{x})]$ and $\mathrm{D}_{\text{\tiny z}}[{\cal L}(\mathsf{x})]$ the pole and zero divisors of ${\cal L}(\mathsf{x})$ on $\mathbb{CP}^{1}$. Given (4) and (5), we will determine $\rho(\mathsf{x})$ from the requirement
$\mathrm{D}_{\text{\tiny p}}[{\cal L}(\mathsf{x})]=\mathrm{D}_{\text{\tiny p}}[{\cal L}^{U}(\rho(\mathsf{x}))],\quad\mathrm{D}_{\text{\tiny z}}[{\cal L}(\mathsf{x})]=\mathrm{D}_{\text{\tiny z}}[{\cal L}^{U}(\rho(\mathsf{x}))].$ (6)
[…]
To determine $U$, one then scans for those gauges under which the zero divisor changes. Typically useful ones are those where the sigma-model fields are collected in a group element, $U=g(\tau,\sigma)\in G$. One can also include constant matrices $U=w\not\in G$ (cf. earlier), but these do not affect the divisor structure.
Furthermore, we consider $\rho:\mathbb{CP}^{1}\rightarrow\mathbb{CP}^{1}$ to be a reflection map that accompanies the spatial reflection, and hence it is an involutive Möbius transformation:
$\rho(\mathsf{x})=\frac{a\mathsf{x}+b}{c\mathsf{x}-a},\quad a^{2}+bc\neq 0,$ (7)
with $\rho^{2}(\mathsf{x})=\mathsf{x}$ and $a,b,c\in\mathbb{C}$ to be determined. It can therefore only permute the marked points of (6) in pairs (or fix them), while preserving their analytic type. For a meromorphic Lax family with finitely many marked points under different gauges, this then leaves only a finite set of involutive $\rho(\mathsf{x})$ and $U$ candidates, that together with different $K_{\sigma_{\star}}$-solutions of (4) can lead to distinct integrable boundary conditions.
II Mixed-flux $AdS_{3}\times S^{3}$
We now apply our approach to boundary integrability in the mixed-flux $AdS_{3}\times S^{3}$ coset sigma-model [14]. The relevant ingredients are reviewed in App. A and B, to which we refer the reader for conventions and notation.
For simplicity, we restrict to the bosonic sector ($g_{\text{\tiny F}}=\tilde{g}_{\text{\tiny F}}=\mathbb{1}$) and fix the coset-gauge $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$, so that (35) applies. There are then two interesting Lax representatives, with $U=\mathbb{1}$ and $U=g_{\text{\tiny B}}$, which reduce to ${\cal L}(\mathsf{x})=f_{1}(\mathsf{x})\mathcal{J}^{\text{\tiny B}}+f_{2}(\mathsf{x})\star\mathcal{J}^{\text{\tiny B}}$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})=(f_{1}(\mathsf{x})-1)\mathcal{R}^{\text{\tiny B}}+f_{2}(\mathsf{x})\star\mathcal{R}^{\text{\tiny B}}$ with $\mathcal{R}^{\text{\tiny B}}=\mathrm{Ad}_{g_{\text{\tiny B}}}\mathcal{J}^{\text{\tiny B}}=g_{\text{\tiny B}}\mathcal{J}^{\text{\tiny B}}g_{\text{\tiny B}}^{-1}=\mathrm{d}g_{\text{\tiny B}}g_{\text{\tiny B}}^{-1}$ and
$f_{1}(\mathsf{x})$ $=\frac{q_{\text{\tiny R}}}{\mathsf{x}^{2}-1}+\frac{q_{\text{\tiny R}}+1}{2},~~f_{2}(\mathsf{x})=\frac{\mathsf{x}q_{\text{\tiny R}}}{1-\mathsf{x}^{2}}+\frac{q_{\text{\tiny NS}}}{2}.$ (8)
Here $q_{\text{\tiny R}},q_{\text{\tiny NS}}\in\mathbb{R}$ denote the amount of RR and NSNS flux respectively, with $q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1$.
Marked points, spectral reflections and boundary conditions.
Both ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ have the same pole divisor: two simple poles at $\mathsf{x}_{\pm}=\pm 1$. In addition, they each have a single, but different, simple zero. In the above coset-gauge, ${\cal L}(\mathsf{x}_{0})=0$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x}_{0}^{g})=0$ with
$\mathsf{x}_{0}=\frac{q_{\text{\tiny R}}-1}{q_{\text{\tiny NS}}},\qquad\mathsf{x}_{0}^{g}=\frac{q_{\text{\tiny R}}+1}{q_{\text{\tiny NS}}}.$ (9)
Notice that when keeping the coset-gauge for $\tilde{g}_{\text{\tiny B}}$ general, ${\cal L}(\mathsf{x})$ would not have a zero, while ${\cal L}^{g}(\mathsf{x})$ remains zero at $\mathsf{x}=\mathsf{x}_{0}^{g}$ (which reduces to $\mathsf{x}=\infty$ in the pure RR case). This corresponds to the standard set-up for coset sigma-models where one expands the Lax representative ${\cal L}^{g}(\mathsf{x})$ to obtain local charges. The possibility of having an additional zero when using $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$ is however special to the $AdS_{3}\times S^{3}$ (permutation-)coset case and is important for its integrable boundary structure.
Given (6), there are then four possible reflections $\rho_{i}(\mathsf{x})$, depending on whether the poles and zeros are fixed or permuted. The behaviour of the zeros further determines the gauges for the Lax connections: when $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ we must take $U=\mathbb{1}$, while for $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}^{g}$ we must take $U=g_{\text{\tiny B}}$. [^5] Note that $\rho(\mathsf{x}_{0}^{g})=\mathsf{x}_{0}^{g}$ with both Lax representatives in the gauge $U=g_{\text{\tiny B}}$ is equivalent to $\rho(\mathsf{x}_{0})=\mathsf{x}_{0},U=\mathbb{1}$ up to redefining $K_{\sigma_{\star}}$. In table 1, we summarise these possibilities, the corresponding solutions of (7) and boundary conditions from (4) for trivial $K_{\sigma_{\star}}$ matrices.
$i$ Zeros Poles $\rho_{i}(\mathsf{x})$ minimal (4)
1 $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\mp}$ $\tfrac{-\mathsf{x}-q_{\text{\tiny NS}}}{q_{\text{\tiny NS}}\mathsf{x}+1},~q_{\text{\tiny NS}}\neq 1$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho(\mathsf{x}))$
2 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\mp}$ $\mathsf{x}^{-1}$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}^{g_{\text{\tiny B}}}_{\tau}(\rho(\mathsf{x}))$
3 $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$ $\mathsf{x}^{-1},~~q_{\text{\tiny R}}=0$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho(\mathsf{x}))$
4 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$
Table 1: The inequivalent Möbius involutions $\rho(\mathsf{x})$.
Specifically, one finds that there is no solution $\rho_{4}(\mathsf{x})$ fixing the poles and permuting the zeros. Yet both poles and zeros are fixed by $\rho_{3}(\mathsf{x})=\mathsf{x}^{-1}$ when $q_{\text{\tiny R}}=0$. However, in this case the Lax connection does not depend on $\mathsf{x}$ and for trivial $K_{\sigma_{\star}}$ one has $T_{\text{\tiny b}}(\mathsf{x})=\mathbb{1}$. The boundary conditions (4) would be trivially satisfied and thus can not provide solutions of the sigma-model boundary equations of motion (LABEL:eq:sm-bcs) (see also [11]).
The interesting cases arise when poles are permuted. Firstly, we have
$\rho_{1}(\mathsf{x})=-\frac{\mathsf{x}+q_{\text{\tiny NS}}}{q_{\text{\tiny NS}}\mathsf{x}+1},\qquad q_{\text{\tiny NS}}\neq 1,$ (10)
which fixes $\mathsf{x}_{0}$. This reduces to the standard reflection $\rho_{1}(\mathsf{x})=-\mathsf{x}$ in the pure RR case ($q_{\text{\tiny NS}}=0$). Mapping residues of the minimal boundary conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho_{1}(\mathsf{x}))$ then in general gives
$\mathcal{J}_{+}^{\text{\tiny B}}\overset{\text{\tiny b}}{=}\frac{1-q_{\text{\tiny NS}}}{1+q_{\text{\tiny NS}}}\mathcal{J}^{\text{\tiny B}}_{-}~~\Rightarrow~~\partial_{\sigma}X^{\mu}\overset{\text{\tiny b}}{=}-q_{\text{\tiny NS}}\partial_{\tau}X^{\mu},$ (11)
[…]
$\partial_{\sigma}\tilde{X}^{\mu_{+}}\overset{\text{\tiny b}}{=}0,\qquad\partial_{\tau}\tilde{X}^{\mu_{-}}\overset{\text{\tiny b}}{=}0,$ (12)
where $\mu_{\pm}$ has cardinality $m_{\pm}$. For the automorphism given in (39), one finds e.g. $m_{+}=4$ and then this corresponds to a D-brane with $p=3$ and ${\cal F}=0$. Nontrivial worldvolume fluxes are typically introduced by dynamical and $\mathsf{x}$-dependent $K_{\sigma_{\star}}$-matrices, as happens for several interesting D-branes in $AdS_{5}\times S^{5}$ and $AdS_{4}\times\mathbb{CP}^{3}$ [30, *Linardopoulos:2022wol, *Demjaha:2025axy, 33]. In the future, it would be interesting to compare with the $\mathsf{x}$-dependent quantum $K_{\sigma_{\star}}$-matrices of [34, *Bielli:2024xuv, *Bielli:2025abu] known for pure RR flux. For generic mixed-flux with (10), an interesting question is also whether compatibility with (LABEL:eq:sm-bcs) requires dynamical reflection data $K_{\sigma_{\star}}$.
The second option permutes both poles and zeros
$\rho_{2}(\mathsf{x})=-\mathsf{x}^{-1},$ (13)
with no restriction on $q_{\text{\tiny NS}},q_{\text{\tiny R}}$. Including automorphisms $W$ and mapping residues, the minimal conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}W\left[{\cal L}^{g_{\text{\tiny B}}}_{\tau}(\rho_{2}(\mathsf{x}))\right]$ now simplify for $q_{\text{\tiny R}}\neq 0$ to
$\mathcal{J}^{\text{\tiny B}}_{\pm}\overset{\text{\tiny b}}{=}-W\left[\mathcal{R}^{\text{\tiny B}}_{\mp}\right]=-(W\mathrm{Ad}_{g_{\text{\tiny B}}})\left[\mathcal{J}^{\text{\tiny B}}_{\mp}\right],$ (14)
implying the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$. Interestingly, for $q_{\text{\tiny R}}=0$, where ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ become $\mathsf{x}$-independent, the boundary conditions enhance to $\mathcal{J}^{\text{\tiny B}}_{+}=-W\left[\mathcal{R}_{-}^{\text{\tiny B}}\right]$ without further constraints. These are precisely the conformal boundary conditions of WZW models which correspond to D-branes wrapping twisted conjugacy classes in $SU(1,1)\times SU(2)$. For $W=\mathbb{1}$ this yields unstable $dS_{2}\times S^{2}$ branes, which we will exclude in the following, while including the nontrivial outer automorphism (40) with $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ yields stable $AdS_{2}\times S^{2}$ branes [25]. In terms of the fields $X^{\mu}=\{\psi,\omega,\tau;\alpha,\beta,\gamma\}$ given in (36), the latter boundary conditions read
$\partial_{\tau}\psi$ $\overset{\text{\tiny b}}{=}0,\qquad$ $\partial_{\sigma}\omega$ $\overset{\text{\tiny b}}{=}\tanh\psi\cosh\omega\partial_{\tau}t,$ (15)
$\partial_{\sigma}t$ $\overset{\text{\tiny b}}{=}\tanh\psi\operatorname{sech}\omega\partial_{\tau}\omega,$
$\partial_{\tau}\alpha$ $\overset{\text{\tiny b}}{=}0,\qquad$ $\partial_{\sigma}\beta$ $\overset{\text{\tiny b}}{=}-\cot\alpha\sin\beta\partial_{\tau}\gamma,$
$\partial_{\sigma}\gamma$ $\overset{\text{\tiny b}}{=}\cot\alpha\csc\beta\partial_{\tau}\beta.$
The conjugacy classes are thus fixed at $\{\psi;\alpha\}$ and extend along $X^{m}=\{t,\omega;\beta,\gamma\}$. With (LABEL:eq:sm-bcs) and (37) the worldvolume flux can then be extracted from (15) as
${\cal F}_{mn}=\tfrac{\sinh 2\psi\cosh\omega}{2}\mathrm{d}\omega\wedge\mathrm{d}t+\tfrac{\sin 2\alpha\sin\beta}{2}\mathrm{d}\gamma\wedge\mathrm{d}\beta,$ (16)
For mixed-flux, however, the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$ fixes the D-brane conjugacy-classes on specific loci, i.e. at
$\psi=0,\qquad\alpha=\{0,\tfrac{\pi}{2},\pi\},$ (17)
on which the worldvolume flux becomes trivial, ${\cal F}=0$. We next show that this restriction can be lifted by allowing nontrivial reflection data (i.e. beyond $K_{\sigma_{\star}}=w_{\text{\tiny o}}$).
Before doing so, we remark that interestingly $\rho_{1}(\mathsf{x})$ at $q_{\text{\tiny NS}}=0$ can equally be interpreted as demanding the $\sigma\rightarrow 2\pi-\sigma$ parity-invariance of ${\cal L}(\mathsf{x})$ after reflection, while $\rho_{2}(\mathsf{x})$ combines requiring invariance under both parity and group element inversion $g_{\text{\tiny B}}\rightarrow g_{\text{\tiny B}}^{-1}$. It would be interesting to relate this to crossing properties of [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda].
Extension to all twisted conjugacy classes.
To restore the full family of stable $AdS_{2}\times S^{2}$ twisted conjugacy classes with two-form (16) at generic $(\psi,\alpha)$, we allow for appropriate matrices $K_{\sigma_{\star}}(\mathsf{x})$ while keeping the same reflection $\rho_{2}(\mathsf{x})=-\mathsf{x}^{-1}$. For this purpose we reverse the logic and use the WZW boundary conditions (15) as input to (4) to determine the required $K_{\sigma_{\star}}(\mathsf{x})$.
We find that it is sufficient to consider the ansatz
$K_{\sigma_{\star}}(\tau;\mathsf{x})=\left(\mathbb{1}+\sum_{k=1}^{6}u_{k}(X^{\mu};\mathsf{x})T_{k}\right)w_{\text{\tiny o}}\ ,$ (18)
with $T_{k}$ the $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ generators given in (27) and $u_{k}$ depending on $\tau$ only through the target-space fields $X^{\mu}=\{\psi,\omega,t;\alpha,\beta,\gamma\}$. Projecting the resulting system (4) onto the generators yields six coupled first-order PDEs in the worldsheet variables $(\tau,\sigma)$. Using (15), we rewrite them in terms of $\partial_{\sigma}\{\psi,\alpha\}$ and $\partial_{\tau}\{t,\omega,\beta,\gamma\}$ only. Imposing that $u_{k}$ does not depend on $\partial_{\sigma}\{\psi,\alpha\}$ (i.e. they can depend on $\psi,\alpha$ but not their derivatives), gives a set of algebraic conditions solved by
$\tfrac{u_{2}(\tau;\mathsf{x})}{u_{1}(\tau;\mathsf{x})}=\cot t,\quad$ $\tfrac{u_{3}(\tau;\mathsf{x})}{u_{1}(\tau;\mathsf{x})}=\tanh\omega\csc t,$ (19)
$\tfrac{u_{5}(\tau;\mathsf{x})}{u_{4}(\tau;\mathsf{x})}=-\cot\gamma,\qquad$ $\tfrac{u_{6}(\tau;\mathsf{x})}{u_{4}(\tau;\mathsf{x})}=-\cot\beta\csc\gamma.$
While this reduces the unfixed functions to $u_{1}$ and $u_{4}$ only, we remarkably find that the resulting six PDEs can still be solved consistently. We simply obtain
$u_{1}=-\sqrt{2}\frac{\sinh 2\psi\cosh\omega\sin t}{z(\mathsf{x})+\cosh 2\psi},~~u_{4}=\frac{\sqrt{2}}{i}\frac{\sin 2\alpha\sin\beta\sin\gamma}{z(\mathsf{x})-\cos 2\alpha},$ (20)
with the rational function $z(\mathsf{x})$ defined as
$z(\mathsf{x})=\frac{(\mathsf{x}-1)(q_{\text{\tiny NS}}(\mathsf{x}+1)+q_{\text{\tiny R}}(\mathsf{x}-1))}{(\mathsf{x}+1)((1-q_{\text{\tiny NS}})(\mathsf{x}-1)+q_{\text{\tiny R}}(\mathsf{x}+1))}.$ (21)
When the denominator of $z(\mathsf{x})$ vanishes, all $u_{k}=0$ and $K_{\sigma_{\star}}=w_{\text{\tiny o}}$. This occurs at $\mathsf{x}=\mathsf{x}_{0}$ and at the WZW point. Moreover also on the loci (17) we get $K_{\sigma_{\star}}=w_{\text{\tiny o}}$. This is thus entirely consistent with the previous subsection. Away from these special points, we thus find that the brane embedding can remain that of a twisted conjugacy class as in boundary WZW models, while $K_{\sigma_{\star}}(\mathsf{x})$ becomes nontrivial showing that the mixed-flux deformation is encoded entirely in the reflection data.
We will report on the extension to the fermionic sector, including the analysis of residual symmetries from $T_{\text{\tiny b}}(\mathsf{x})$ and the open-string spectrum, elsewhere.
Acknowledgements.
We thank D. Bielli, C. Kristjansen, S. Lacroix, G. Linardopoulos, B. Nairz, D. Polvara, A. Retore, K. Zarembo for useful discussions, and A. Retore, K. Zarembo and especially S. Lacroix for comments on the draft. This work is partly based on the Master thesis of JCG prepared at ETH Zürich and we thank N. Beisert for acting as co-advisor. JCG’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”. SD is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and the NCCR SwissMAP. SD also thanks the organisers and participants of the workshop Higher-$d$ integrability (Favignana, 2025) where part of this work was presented.
Appendix A Sigma-model boundary conditions
We formulate integrable boundary conditions through a relation on the Lax connection (4), which guarantees conservation of charges generated by $T_{\text{\tiny b}}(\mathsf{x})$. At the level of the sigma-model, this must translate into boundary conditions on the fields $X^{\mu}$, $\mu=0,\ldots,D-1$, that are compatible with the variational principle.
The bosonic action of a standard two-derivative sigma-model on $\Sigma=\mathbb{R}_{\tau}\times[0,\pi]_{\sigma}$ can be written as  [^6] Here we use the flat sigma-model metric $\eta_{\alpha\beta}=\mathrm{diag}(1,-1)$, with $\sigma^{\alpha}=(\tau,\sigma)$, and $\epsilon_{\tau\sigma}=1$. In general the target space fields (metric and B-field) are curved, $G_{\mu\nu}:=G_{\mu\nu}(X)$ and $B_{\mu\nu}:=B_{\mu\nu}(X)$ and $T$ is the sigma-model coupling (i.e. string tension). As an action of open strings (22) is written in conformal gauge.
$S={}$ $\frac{T}{2}\int_{\Sigma}\mathrm{d}^{2}\sigma\;\partial_{\alpha}X^{\mu}\big(\eta^{\alpha\beta}G_{\mu\nu}+\epsilon^{\alpha\beta}B_{\mu\nu}\big)\partial_{\beta}X^{\nu}$ (22)
$+\int_{\partial\Sigma}\mathrm{d}\tau\;A_{m}\,\partial_{\tau}X^{m},$
where $m,n=0,\cdots,p\leq D-1$ are the directions along which $X^{\mu}|_{\partial\Sigma}$ is allowed to fluctuate, i.e. (for open strings) those along a Dpp-brane with a $U(1)$ gauge field $A_{m}:=A_{m}(X)$ and field strength $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$. The boundary equations of motion are then solved by
Dirichlet: $\partial_{\tau}X^{\hat{m}}\overset{\text{\tiny b}}{=}0,$ (23)
gen. Neumann: $G_{mn}\partial_{\sigma}X^{n}\overset{\text{\tiny b}}{=}{\cal F}_{mn}\partial_{\tau}X^{n},$
with $\hat{m}=p+1,\cdots,D-1$ and ${\cal F}_{mn}(X)=B_{mn}(X)+T^{-1}F_{mn}(X)$ the gauge-invariant worldvolume flux. Here we assumed adapted coordinates in which the target-space fields split orthogonally in the Neumann ($m$) and Dirichlet directions ($\hat{m}$). Note that the conditions (LABEL:eq:sm-bcs) are rather restrictive, given the symmetry properties of each term in the generalised Neumann conditions.
The problem addressed in this Letter can be restated as determining those Dirichlet and Neumann conditions which are guaranteed to preserve classical integrability. For the sigma-models considered here, the Lax connection ${\cal L}(\mathsf{x})\in\mathbb{C}(\mathsf{x})\otimes\mathfrak{g}$ is a linear combination of (projections of) $\mathfrak{g}$-valued one-form currents $\cal J$} and $\star{\cal J}$ with $\mathsf{x}$-dependent coefficients. The currents themselves are linear in $\mathrm{d}X^{\mu}$ so that one can introduce a frame field $E$ as
$\mathcal{J}_{\alpha}=E_{\mu}{}^{A}(X)T_{A}\partial_{\alpha}X^{\mu},\quad(\star\mathcal{J})_{\alpha}=\mathcal{J}_{\beta}\epsilon^{\beta}{}_{\alpha}$, (24)
with $T_{A}$ the generators of $\mathfrak{g}$. Imposing (4) for chosen $(\rho,U,K_{\sigma_{\star}})$ therefore yields linear gluing relations between $\partial_{\sigma}X^{\mu}$ and $\partial_{\tau}X^{\mu}$ at $\partial\Sigma$ which must be matched to the expressions (LABEL:eq:sm-bcs) for consistency with the variational principle, and hence determine the dimension $p$ and the flux $\mathcal{F}_{mn}$ (equivalently $F_{mn}$ for given $B_{mn}$). Likewise, one may start from local D-brane data $(p,{\cal F})$, and solve (4) for $(\rho,U,K_{\sigma_{\star}})$ using the relations (LABEL:eq:sm-bcs).
Appendix B Mixed-flux $AdS_{3}\times S^{3}$ bulk sigma-model
In this appendix, we collect the ingredients for the boundary integrability analysis of the mixed NSNS/RR flux string on $AdS_{3}\times S^{3}\times M_{4}$, with $M_{4}$ a four-dimensional compact manifold required to complete the type IIB supergravity background (see [12, *Seibold:2024qkh] for reviews). Since the integrable structure relevant for our purposes is encoded in the $AdS_{3}\times S^{3}$ sector, we will restrict attention to its Green-Schwarz formulation as a $\mathbb{Z}_{4}$-graded supercoset sigma-model [14] (see also [37, *Hoare:2014oua]), while keeping $M_{4}$ implicit throughout.
Supercoset.
Concretely, the worldsheet degrees of freedom in this sector are described by $\widehat{F}/F^{(0)}$ with
$\widehat{F}=G\times G,\qquad G=PSU(1,1|2),$ (25)
and $F^{(0)}$ the bosonic diagonal subgroup of $\widehat{F}$, i.e. $(h,h)\in F^{(0)}$ for $h\in SU(1,1)\times SU(2)$. The Lie superalgebra $\hat{\mathfrak{f}}=\mathfrak{g}\oplus\mathfrak{g}$, with $\mathfrak{g}=\mathfrak{psu}(1,1|2)$, of $\widehat{F}$ admits a $\mathbb{Z}_{4}$ grading
$\hat{\mathfrak{f}}=\mathfrak{f}^{(0)}\oplus\mathfrak{f}^{(1)}\oplus\mathfrak{f}^{(2)}\oplus\mathfrak{f}^{(3)},$ (26)
with $\mathfrak{f}^{(0)}$ the Lie algebra corresponding to $F^{(0)}$. Here $\mathfrak{g}=\mathfrak{psu}(1,1|2)$, whose bosonic subalgebra $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ is generated by
$\begin{gathered}T_{1,2}=\tfrac{1}{\sqrt{2}}\left(\sigma_{1,3}\oplus 0\right),\qquad T_{3}=\tfrac{i}{\sqrt{2}}\left(\sigma_{2}\oplus 0\right),\\ T_{4,5,6}=\tfrac{1}{\sqrt{2}}\left(0\oplus\sigma_{1,2,3}\right),\end{gathered}$ (27)
with $\sigma_{1,2,3}$ the Pauli matrices. For a complete superalgebra realisation of $\mathfrak{g}$} and $\hat{\mathfrak{f}}$, see e.g. App. C of [39].
[…]
$\mathfrak{J}^{(0)}$ $=\begin{pmatrix}\mathcal{J}^{(0)}&\\ &\mathcal{J}^{(0)}\end{pmatrix}$$\quad$ $\mathfrak{J}^{(1)}=\begin{pmatrix}\mathcal{J}^{(1)}&\\ &-i\mathcal{J}^{(1)}\end{pmatrix}$, (28)
$\mathfrak{J}^{(2)}$ $=\begin{pmatrix}\mathcal{J}^{(2)}&\\ &-\mathcal{J}^{(2)}\end{pmatrix}$$\quad$ $\mathfrak{J}^{(3)}=\begin{pmatrix}\mathcal{J}^{(3)}&\\ &i\mathcal{J}^{(3)}\end{pmatrix}$,
[…]
$\operatorname{STr}(\mathfrak{J})=\mathrm{Str}(\mathcal{J})+\mathrm{Str}(\widetilde{\mathcal{J}}),\quad\widetilde{\operatorname{STr}}(\mathfrak{J})=\operatorname{STr}(\mathfrak{W}\mathfrak{J}),$ (29)
with $\mathrm{Str}$ the supertrace of $\mathfrak{g}$ and $\mathfrak{W}=\mathrm{diag}(\mathbb{1},-\mathbb{1})$.
Action and integrability.
The bulk sigma-model action with mixed-flux contributions is [14, 37, *Hoare:2014oua]
$S_{\text{\tiny cl.}}={}$ $\frac{1}{2}\int_{\Sigma}\operatorname{STr}\left(\mathfrak{J}^{(2)}\wedge\star\mathfrak{J}^{(2)}+q_{\text{\tiny R}}\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(3)}\right)$ (30)
$+q_{\text{\tiny NS}}\int_{\cal B}\widetilde{\operatorname{STr}}\left(\frac{2}{3}\mathfrak{J}^{(2)}\wedge\mathfrak{J}^{(2)}\wedge\mathfrak{J}^{(2)}\right$.
$+\left.\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(3)}\wedge\mathfrak{J}^{(2)}+\mathfrak{J}^{(3)}\wedge\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(2)}\right),$
[…]
Classical integrability, $\kappa$-symmetry and conformal invariance to one-loop in $\alpha^{\prime}$ require the following relation between the flux parameters [14]
$q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1.$ (31)
The flat Lax connection then takes the form
${\cal L}(\mathsf{x})={}$ ${\cal J}^{(0)}+q_{\text{\tiny R}}\frac{\mathsf{x}^{2}+1}{\mathsf{x}^{2}-1}\mathcal{J}^{(2)}+\left(q_{\text{\tiny NS}}-\frac{2q_{\text{\tiny R}}\mathsf{x}}{\mathsf{x}^{2}-1}\right)\star\mathcal{J}^{(2)}$
$+\left(\mathsf{x}+\frac{q_{\text{\tiny R}}}{1-q_{\text{\tiny NS}}}\right)\sqrt{\frac{q_{\text{\tiny R}}(1-q_{\text{\tiny NS}})}{\mathsf{x}^{2}-1}}\mathcal{J}^{(1)}$ (32)
$+\left(\mathsf{x}-\frac{q_{\text{\tiny R}}}{1+q_{\text{\tiny NS}}}\right)\sqrt{\frac{q_{\text{\tiny R}}(1+q_{\text{\tiny NS}})}{\mathsf{x}^{2}-1}}\mathcal{J}^{(3)}.$
At the pure RR point $q_{\text{\tiny R}}=1$, $q_{\text{\tiny NS}}=0$ this Lax corresponds to [40], while at the pure NSNS/WZW point it degenerates and does not evidently capture the full dynamics. In that case one can instead work with holomorphic current algebras rather than Lax connections.
Parametrisation and background.
The sigma-model (22) is invariant under local, right-acting, $F^{(0)}$ gauge transformations. We can therefore take
$g=g_{\text{\tiny F}}g_{\text{\tiny B}},\qquad\tilde{g}=\tilde{g}_{\text{\tiny F}}\tilde{g}_{\text{\tiny B}}=\tilde{g}_{\text{\tiny F}},$ (33)
[…]
$\mathcal{J}=\mathcal{J}^{\text{\tiny B}}+g_{\text{\tiny B}}^{-1}\mathcal{J}^{\text{\tiny F}}g_{\text{\tiny B}},\qquad\widetilde{\mathcal{J}}=\widetilde{\mathcal{J}}^{\text{\tiny F}},$ (34)
with $\mathcal{J}^{\text{\tiny B}}=g_{\text{\tiny B}}^{-1}\mathrm{d}g_{\text{\tiny B}}\in\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$, $\mathcal{J}^{\text{\tiny F}}=g_{\text{\tiny F}}^{-1}\mathrm{d}g_{\text{\tiny F}}\in\mathfrak{g}$ and $\widetilde{\mathcal{J}}^{\text{\tiny F}}=\tilde{g}_{\text{\tiny F}}^{-1}\mathrm{d}\tilde{g}_{\text{\tiny F}}\in\mathfrak{g}$. In the main text, we furthermore focus on the bosonic sector of $\mathfrak{psu}(1,1|2)$ and thus take $g_{\text{\tiny F}},\tilde{g}_{\text{\tiny F}}=\mathbb{1}$. Altogether, this implies the simplifications
$\left.\mathcal{J}^{(0)}\right|_{g_{\text{\tiny B}}}=\left.\mathcal{J}^{(2)}\right|_{g_{\text{\tiny B}}}=\frac{1}{2}\mathcal{J}^{\text{\tiny B}},\qquad\left.\mathcal{J}^{(1,3)}\right|_{g_{\text{\tiny B}}}=0.$ (35)
We will then parametrise the bosonic element $g_{B}=\mathrm{diag}(g_{SU(1,1)}~|~g_{SU(2)})$ as
$g_{SU(1,1)}$ $=\sinh\psi~\sigma_{3}+\cosh\psi~e^{\tfrac{it}{2}\sigma_{2}}e^{\omega\sigma_{1}}e^{\tfrac{it}{2}\sigma_{2}},$ (36)
$g_{SU(2)}$ $=e^{-\tfrac{i\gamma}{2}\sigma_{3}}e^{\tfrac{i\beta}{2}\sigma_{1}}e^{i\alpha\sigma_{3}}e^{-\tfrac{i\beta}{2}\sigma_{1}}e^{\tfrac{i\gamma}{2}\sigma_{3}},$
where ${\psi,\omega,t}\in\mathbb{R}^{3}$ are global coordinates on $AdS_{3}$, with $\{\omega,t\}$ parametrising $AdS_{2}$ fibred over $\psi$, while the coordinates $\beta\in[0,\pi)$ and $\gamma\in[0,2\pi)$ parametrise an $S^{2}$ fibered over the interval $\alpha\in[0,\pi)$, forming the three-sphere $S^{3}$. The metric and NSNS three-form $H=\mathrm{d}B$ can then be expressed as
$\mathrm{d}s^{2}={}$ $\mathrm{d}\psi^{2}+\cosh^{2}\psi\left(-\cosh^{2}\omega\mathrm{d}t^{2}+\mathrm{d}\omega^{2}\right)$ (37)
$+\mathrm{d}\alpha^{2}+\sin^{2}\alpha\left(\mathrm{d}\beta^{2}+\sin^{2}\beta\mathrm{d}\gamma^{2}\right),$
$H={}$ $2q_{\text{\tiny NS}}\cosh^{2}\psi\cosh\omega\mathrm{d}\psi\wedge\mathrm{d}t\wedge\mathrm{d}\omega$
$+2q_{\text{\tiny NS}}\sin^{2}\alpha\sin\beta\mathrm{d}\alpha\wedge\mathrm{d}\beta\wedge\mathrm{d}\gamma.$
When including the fermionic sector, the RR 3-form flux is $F_{3}=e^{-\Phi_{0}}q_{\text{\tiny R}}(H/q_{\text{\tiny NS}})$ for constant dilaton $\Phi(X)=\Phi_{0}$.
Outer automorphisms of $\bm{SU(1,1)\times SU(2)}$.
Automorphisms $W$ of $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ which preserve the ad-invariant metric form the group $SO(2,1)\times SO(3)$. Explicit matrix realisations can be defined as $W(y)=wyw^{-1}=y^{A}W_{A}{}^{B}T_{B}$ for $y=y^{A}T_{A}\in\mathfrak{g}$. The elements relevant to the boundary integrability discussion of the main text are those $W$ which are involutive, $W^{2}=\mathbb{1}$. Up to conjugation (and combinations), the ones that do not mix $\mathfrak{su}(1,1)$ and $\mathfrak{su}(2)$ and that (i) are inner are
$W_{\text{\tiny inn}}=\left\{\begin{array}[]{ll}\mathbb{1}_{3}\oplus\mathbb{1}_{3}\\ \mathrm{diag}(-1,-1,1)\oplus\mathbb{1}_{3}\\ \mathbb{1}_{3}\oplus\mathrm{diag}(-1,-1,1)\end{array}\right.~\in SO^{+}(2,1)\times SO(3),$ (38)
or (ii) are outer are
$W_{\text{\tiny out.}}=\mathrm{diag}(-1,1,-1)\oplus\mathbb{1}_{3}\in SO^{-}(2,1)\times SO(3).$ (39)
The latter is realised as $W_{\text{\tiny out.}}(X)=w_{\text{\tiny o}}Xw_{\text{\tiny o}}^{-1}$ with
$w_{\text{\tiny o}}=\mathrm{diag}(1,-1)\oplus\mathbb{1}_{2}\not\in SU(1,1)\times SU(2).$ (40)