### v2

# Analytic approach to boundary integrability with application to mixed-flux $AdS_{3}\times S^{3}$

Abstract
Boundary integrability provides rare analytic control over field theories with interfaces, from quantum impurity problems to open string dynamics. We propose an analytic approach for integrable boundaries in two-dimensional sigma-models that determines admissible reflection maps directly from the divisor structure of the Lax connection. Applied to open strings on $AdS_{3}\times S^{3}$ with mixed NSNS-RR flux, we find two branches of integrable boundaries: one restricted to pure RR flux, and another admitting D-branes wrapping twisted conjugacy classes for generic flux. At the WZW point, these reduce to the known conformal D-branes, opening a path to comparison with conformal perturbation theory. More broadly, our framework suggests generalisations of standard lattice constructions that may enlarge existing classifications of integrable boundaries.
preprint: HU-EP-26/02-RTG
Field theories with interfaces play an important role in both high-energy and condensed-matter physics. In open string theory, worldsheet boundary conditions describe D-branes, while in condensed-matter and lattice systems they model boundary critical phenomena and impurities such as the Kondo effect. When the bulk theory is integrable, the situation becomes especially interesting, as an infinite set of conserved charges provides a rare source of analytic control for interacting dynamics. The presence of an interface probes this structure in a nontrivial way. Geometric symmetries are typically broken, and it is far from obvious that an infinite tower of conserved charges can survive. When they do, notions of boundary integrability have e.g. opened access to exact reflection data [1] and methods for computing defect observables in AdS/CFT [2]. Identifying which physical interfaces are compatible with integrable dynamics is therefore a basic structural question. In this Letter, we address it for two-dimensional sigma-models using simple analytical arguments from a novel perspective.
The starting point is an integrable sigma-model on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$ with an on-shell flat Lax connection ${\cal L}(\mathsf{x})$, [^1] For brevity, and where possible, we will suppress explicit time ($\tau\in\mathbb{R}_{\tau}:=\mathbb{R}$) or space ($\sigma\in S^{1}_{\sigma}:=S^{1}$ with radius 1) dependence. i.e. $\mathrm{d}{\cal L}(\mathsf{x})+{\cal L}(\mathsf{x})\wedge{\cal L}(\mathsf{x})=0$, that in general is meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$ and represents the equations of motion for every value of the spectral parameter $\mathsf{x}$. In this case, and under periodic boundary conditions ${\cal L}(\sigma+2\pi;\mathsf{x})={\cal L}(\sigma;\mathsf{x})$, the monodromy matrix $T_{\cal L}(2\pi,0;\mathsf{x})$, with
$T_{\cal L}(b,a;\mathsf{x})=\overset{\leftarrow}{\mathcal{P}}\exp\left(-\int^{b}_{a}d\sigma{\cal L}_{\sigma}(\mathsf{x})\right),~~~~a,b\in\mathbb{R},$ (1)
is a generator of conserved charges, as its eigenvalues are independent of $\tau$ for all $\mathsf{x}$ (see e.g. the lecture notes [3]).
On an interval $\sigma\in[0,\pi]$, boundary integrability requires a modification of this generator. The prototype is obtained by following Sklyanin’s construction [4, 5], where one forms a double-row monodromy matrix $T_{\text{\tiny b}}(\mathsf{x})$ by gluing $T_{\cal L}(\pi,0;\mathsf{x})$ over the interval $\sigma\in[0,\pi]$ to a reflected copy. On the lattice, Sklyanin’s approach applies to integrable models with crossing-symmetric $R$-matrices of difference form [5]. The reflection compatible with integrability then has two components: a spatial flip reversing the order of sites, and a spectral reflection $\mathsf{x}\mapsto\rho(\mathsf{x})=-\mathsf{x}$ implementing crossing symmetry. Additional boundary degrees of freedom are encoded in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ at the endpoints $\sigma_{\star}=0,\pi$, which must satisfy the Boundary Yang-Baxter Equation (BYBE) to preserve integrability. The analogue in the sigma-model is to apply a worldsheet parity flip $\sigma\rightarrow 2\pi-\sigma$ combined with a spectral reflection $\rho(\mathsf{x})$ that ensures invariance of the Lax connection, see e.g. [4, 6, 7, 8]. One can then effectively set $T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T_{\cal L}^{-1}(\pi,0;\rho(\mathsf{x}))K_{\pi}^{-1}(\mathsf{x})T_{\cal L}(\pi,0;\mathsf{x})$ and solve the conditions for $K_{\sigma_{\star}}$ under which infinite charges are generated by $T_{\text{\tiny b}}(\mathsf{x})$.
However, there are limitations to taking this prescription directly as a definition of boundary integrability for sigma-models. On the lattice, $\rho(\mathsf{x})$ is dictated by the crossing structure of a chosen bulk $R$-matrix. For an integrable sigma-model, such input is often unavailable (or indirect through the sigma-model $S$-matrix), and the defining object is just a Lax connection ${\cal L}(\mathsf{x})$. In addition, there need not exist a $\rho(\mathsf{x})$ that leaves ${\cal L}(\mathsf{x})$ invariant after a parity flip. In these cases, there is no given principle to select an appropriate generator of conserved charges.
In this Letter, we will instead bypass direct input on bulk parity or crossing invariance and determine $\rho(\mathsf{x})$ from the analytic structure of the Lax connection. [^2] For alternative approaches via parity, yet with additional assumptions, see [9, *Corrigan:1996nt]. Here we will take an analytic approach that moreover admits a natural origin in 4-dimensional Chern-Simons theory, as will be reported elsewhere [11]. A central shift in our approach is to fix $\rho(\mathsf{x})$ by treating the reflected copy of ${\cal L}(\mathsf{x})$ modulo gauge transformations that preserve the flatness condition. Our prescription to fix both $\rho(\mathsf{x})$ and the gauge representative $U$ is then to simply demand that the divisor of zeros and poles of ${\cal L}(\mathsf{x})$ on $\mathbb{CP}^{1}$ is preserved under the reflection. Once done, residual boundary reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ can be introduced as above to select distinct integrable interfaces.
As an important case, we apply this approach to the string sigma-model on $AdS_{3}\times S^{3}\times M_{4}$ with mixed NSNS and RR flux; for reviews see [12, *Seibold:2024qkh]. Its classical Lax connection is known for all flux values [14] and illustrates the above limitation: the parity-based criterion alone does not fix $\rho(\mathsf{x})$ at nontrivial NSNS flux. Moreover, while worldsheet methods have led to $S$-matrices [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda], a dual CFT and lattice formulation remains absent for generic fluxes. This makes the mixed-flux model an ideal setting to demonstrate our approach to boundary integrability.
Boundary integrability of the mixed-flux $AdS_{3}\times S^{3}$ background is also inherently motivated by exact worldsheet studies. At the pure NSNS point, the bosonic theory is described by an exactly solvable WZW model on $SU(1,1)\times SU(2)$ [20, *Maldacena:2000hw], in which CFT-preserving D-branes wrap twisted conjugacy classes that admit a boundary CFT (BCFT) description [22, *Stanciu:1999id, *Figueroa-OFarrill:2000gfl, 25]. Turning on RR-flux corresponds to a marginal deformation and, for pointlike branes, its effect on the open-string spectrum has been analysed through boundary conformal perturbation theory, leading to all-orders resummed results, while in the closed-string sector the RR perturbation is singular [26]. The boundary sector thus provides one of the few controlled windows on mixed-flux physics. It was in fact suggested in [14] that identifying integrable boundary conditions could provide further structural information at finite flux and offer a bridge to the BCFT picture. The present Letter realises this at the classical bosonic level. Imposing our analytic criterion on the mixed-flux Lax, we find two branches of integrable boundary conditions: one compatible with the sigma-model action only in the pure RR case, while the other persists for generic flux and includes all D-branes wrapping twisted conjugacy classes. Interestingly, the former arises by identifying $\rho(\mathsf{x})$ from the invariance of ${\cal L}(\mathsf{x})$ under parity alone, whereas the latter requires combining parity with a gauge transformation of the Lax connection. It is in the second branch that we have a rigid class of integrable branes whose embedding coincides with the conformal WZW branes. We show that their worldvolume two-form can remain flux-independent, with the mixed-flux contribution encoded entirely in dynamical reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$. [^3] This is compatible with the $\lambda$-deformed WZW case, where integrable D-branes also remain twisted conjugacy classes but for which the gauge-invariant worldvolume two-form acquires a non-trivial $\lambda$–dependence given the $\lambda$–deformed metric [27].
Outlook. For mixed-flux $AdS_{3}\times S^{3}$, our results thus provide a natural setup to compare integrability-based data with conformal perturbation theory around the solvable WZW point, where the same twisted conjugacy classes underlie the BCFT description. Including the fermionic sector and analysing the open string spectrum would enable comparison with the all-order spectra of [26] and open the way to extending the analysis beyond the pointlike brane sector. We leave this for future work.
More broadly, our approach not only generalises sigma-model constructions but also suggests a possible generalisation for lattice models. Traditionally, the BYBE assumes that both rows in the double-row monodromy share the same $R$-matrix. By allowing different representatives in each row, mirroring different gauge choices of ${\cal L}(\mathsf{x})$, a natural question is whether classifications can be enlarged to include new integrable interfaces that yield distinct states after reflection. Given the scarcity of integrable models with interfaces—for defect observables in holography, benchmarking impurity physics, engineering quantum circuits, etc.—this could provide valuable new ground for applications of exact methods.
I Analytic criteria
To analyse integrable two-dimensional sigma-models with boundaries as outlined in the introduction, we start from bulk data: an on-shell flat Lax connection ${\cal L}(\mathsf{x})$, generally in an $n\times n$ matrix representation of the loop algebra ${\mathbb{C}(\mathsf{x})}\otimes\mathfrak{g}$. Under periodic conditions on $S^{1}_{\sigma}$, towers of charges are then generated by $T_{{\cal L}}(2\pi,0;\mathsf{x})$. One can also allow for flatness-preserving gauge transformations
${\cal L}^{U}(\mathsf{x})=U{\cal L}(\mathsf{x})U^{-1}-\mathrm{d}UU^{-1},$ (2)
with any dynamical $n\times n$ matrix $U:=U(\tau,\sigma;\mathsf{x})$. Then $T_{{\cal L}^{U}}(2\pi,0;\mathsf{x})=U(\tau,2\pi;\mathsf{x})T_{\cal L}(2\pi,0;\mathsf{x})U(\tau,0;\mathsf{x})^{-1}$ is also a generator of charges when $U$ is periodic in $\sigma$.
With open boundaries, we propose the following ansatz of the double-row monodromy matrix
$T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T^{-1}_{{\cal L}^{U}}(\pi,0;\rho(\mathsf{x}))K_{\pi}(\mathsf{x})^{-1}T_{\cal L}(\pi,0;\mathsf{x}),$ (3)
which generates an infinite tower of conserved charges under the integrable boundary conditions [^4] Flatness of the Lax and the boundary conditions (4) indeed imply that $\partial_{\tau}\mathrm{Tr}T_{\text{\tiny b}}(\mathsf{x})^{p}=0$ for any integer $p$, so that expanding the eigenvalues of $T_{\text{\tiny b}}(\mathsf{x})$ in $\mathsf{x}$ produces conserved objects.
${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}K_{\sigma_{\star}}(\mathsf{x}){\cal L}_{\tau}^{U}(\rho(\mathsf{x}))K_{\sigma_{\star}}^{-1}(\mathsf{x})-\partial_{\tau}K_{\sigma_{\star}}(\mathsf{x})K_{\sigma_{\star}}^{-1}(\mathsf{x}),$ (4)
with $\overset{\text{\tiny b}}{=}$” denoting evaluation at the boundary $\sigma_{\star}=0,\pi$ and $K_{\sigma_{\star}}:=K(\tau;\mathsf{x})$ $\sigma$-independent reflection matrices.
Remarks: (i) Constant matrices $K_{\sigma_{\star}}$ or $U=w$, can implement automorphisms of $\mathfrak{g}$ when $W(y)=wyw^{-1}\in\mathfrak{g}$ for $y\in\mathfrak{g}$ [7]. At the lattice level, these can give different representation choices for the reflected bulk sites. In suitable realisations, this covers the common alternative convention where the inverse in (3) is replaced by transposition. (ii) Only the endpoint values $U(\tau,\sigma_{\star})$ enter in (4), and given (2) one may thus see the bulk gauge $U$ as part of a boundary matrix $\tilde{K}=KU$. However, for us the role of $U$ is fundamentally distinct: it will determine relevant $\rho(\mathsf{x})$ for the reflected bulk region, prior to solving for boundary data. In other words, $K_{\sigma_{\star}}$ is the residual gauge freedom at the boundary after reflection. This viewpoint will let us generalise the prescription of [4, 6, 7, 8, 28, 27, 29] beyond parity-invariant gluing.
Fixing $U$ and $\rho(\mathsf{x})$ from bulk analyticity.
Our main point is that $U$ and $\rho(\mathsf{x})$ can be fixed intrinsically from the analytic structure of ${\cal L}(\mathsf{x})$, mirroring the lattice logic where $\rho(\mathsf{x})$ is derived from bulk properties of the $R$-matrix (rather than boundary $K$-matrices).
A first useful criterion is a choice of “physical” representatives for the unreflected and reflected rows, such that $T_{\text{\tiny b}}(\mathsf{x})$ can generate a set of local charges. Concretely, one requires that there exist points $\mathsf{x}_{0},\mathsf{x}^{\prime}_{0}\in\mathbb{CP}^{1}$ where
${\cal L}_{\sigma}(\mathsf{x}_{0})=0,\qquad{\cal L}_{\sigma}^{U}(\mathsf{x}^{\prime}_{0})=0,$ (5)
so that $T_{\text{\tiny b}}(\mathsf{x})$ has a local expansion around $\mathsf{x}_{0}$ when $\mathsf{x}^{\prime}_{0}=\rho(\mathsf{x}_{0})$ for some $\rho(\mathsf{x})$. [^5] Note that $K_{\sigma_{\star}}$ is $\sigma$-independent and thus, when regular at $\mathsf{x}_{0}$, does not alter locality. In practice, relevant representatives are found by scanning for those gauges under which ${\cal L}^{U}_{\sigma}$ remains local while its zero divisor changes. Note, however, that this is a strong requirement: a local $U(\mathsf{x}_{0}^{\prime})$ where ${\cal L}_{\sigma}^{U}(\mathsf{x}^{\prime}_{0})=0$ requires ${\cal L}_{\sigma}(\mathsf{x}_{0}^{\prime})$ to be off-shell flat. Relaxing these criteria would formally enlarge the possibilities, but could either introduce incompatibilities with the conditions (4) for trivial $K_{\sigma_{\star}}$ or obstruct their interpretation as local boundary actions as e.g. in (22).
[…]
$\mathrm{D}_{\text{\tiny p}}[{\cal L}(\mathsf{x})]=\mathrm{D}_{\text{\tiny p}}[{\cal L}^{U}(\rho(\mathsf{x}))],\quad\mathrm{D}_{\text{\tiny z}}[{\cal L}(\mathsf{x})]=\mathrm{D}_{\text{\tiny z}}[{\cal L}^{U}(\rho(\mathsf{x}))].$ (6)
Furthermore, we consider $\rho:\mathbb{CP}^{1}\rightarrow\mathbb{CP}^{1}$ to be a nontrivial reflection map that accompanies the spatial reflection, and hence as the involutive Möbius transformation:
$\rho(\mathsf{x})=\frac{a\mathsf{x}+b}{c\mathsf{x}-a},\quad a^{2}+bc\neq 0,$ (7)
with $\rho^{2}(\mathsf{x})=\mathsf{x}$ and $a,b,c\in\mathbb{C}$ to be determined. It can therefore only permute the marked points of (6) in pairs (or fix them), while preserving their analytic type. For a meromorphic Lax connection with finitely many marked points under different gauges, this then leaves only a finite set of possible $\rho(\mathsf{x})$ and $U$ candidates, that together with different $K_{\sigma_{\star}}$ solutions of (4) can lead to distinct integrable boundary conditions.
II Mixed-flux $AdS_{3}\times S^{3}$
We now apply our approach to boundary integrability in the mixed-flux $AdS_{3}\times S^{3}$ coset sigma-model [14]. The relevant ingredients are reviewed in App. A and B, to which we refer the reader for conventions and notation.
For simplicity, we restrict to the bosonic sector ($g_{\text{\tiny F}}=\tilde{g}_{\text{\tiny F}}=\mathbb{1}$) and fix the coset-gauge $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$, so that (35) applies. There are then two interesting Lax representatives, with $U=\mathbb{1}$ and $U=g_{\text{\tiny B}}$, which reduce to ${\cal L}(\mathsf{x})=f_{1}(\mathsf{x})\mathcal{J}^{\text{\tiny B}}+f_{2}(\mathsf{x})\star\mathcal{J}^{\text{\tiny B}}$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})=(f_{1}(\mathsf{x})-1)\mathcal{R}^{\text{\tiny B}}+f_{2}(\mathsf{x})\star\mathcal{R}^{\text{\tiny B}}$ with $\mathcal{R}^{\text{\tiny B}}=\mathrm{Ad}_{g_{\text{\tiny B}}}\mathcal{J}^{\text{\tiny B}}=g_{\text{\tiny B}}\mathcal{J}^{\text{\tiny B}}g_{\text{\tiny B}}^{-1}=\mathrm{d}g_{\text{\tiny B}}g_{\text{\tiny B}}^{-1}$ and
$f_{1}(\mathsf{x})$ $=\frac{q_{\text{\tiny R}}}{\mathsf{x}^{2}-1}+\frac{q_{\text{\tiny R}}+1}{2},~~f_{2}(\mathsf{x})=\frac{\mathsf{x}q_{\text{\tiny R}}}{1-\mathsf{x}^{2}}+\frac{q_{\text{\tiny NS}}}{2}.$ (8)
Here $q_{\text{\tiny R}},q_{\text{\tiny NS}}\in\mathbb{R}$ denote the amount of RR and NSNS flux of the full theory respectively, related as $q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1$.
Marked points, spectral reflections and boundary conditions.
Both ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ have the same pole divisor: two simple poles at $\mathsf{x}_{\pm}=\pm 1$. In addition, they each have a single, but different, simple zero. In the above coset-gauge, ${\cal L}(\mathsf{x}_{0})=0$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x}_{0}^{g})=0$ with [^7] When keeping the coset-gauge for $\tilde{g}_{\text{\tiny B}}$ general, ${\cal L}(\mathsf{x})$ would have no zero while ${\cal L}^{g}(\mathsf{x})$ remains zero at $\mathsf{x}=\mathsf{x}_{0}^{g}$ (which in the pure RR case reduces to $\mathsf{x}=\infty$). This corresponds to the standard set-up for coset models where one expands ${\cal L}^{g}(\mathsf{x})$ to obtain local charges. The possibility of having an additional zero when using $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$ is however special to $AdS_{3}\times S^{3}$ (permutation-)cosets.
$\mathsf{x}_{0}=\frac{q_{\text{\tiny R}}-1}{q_{\text{\tiny NS}}},\qquad\mathsf{x}_{0}^{g}=\frac{q_{\text{\tiny R}}+1}{q_{\text{\tiny NS}}}.$ (9)
Given (6), there are then four possible reflections $\rho_{i}(\mathsf{x})$, depending on whether the poles and zeros are fixed or permuted. The behaviour of the zeros further determines the necessary gauges for the Lax connections: when $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ we must take $U=\mathbb{1}$, while for $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}^{g}$ we must take $U=g_{\text{\tiny B}}$. [^8] Note that $\rho(\mathsf{x}_{0}^{g})=\mathsf{x}_{0}^{g}$ with both Lax representatives in the gauge $U=g_{\text{\tiny B}}$ is equivalent to $\rho(\mathsf{x}_{0})=\mathsf{x}_{0},U=\mathbb{1}$ up to redefining $K_{\sigma_{\star}}$. In table 1, we summarise the possibilities, the solutions of (7) and the “minimal” boundary conditions from (4), i.e. for trivial $K_{\sigma_{\star}}$ matrices.
$i$ Zeros Poles $\rho_{i}(\mathsf{x})$ minimal (4)
1 $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\mp}$ $\tfrac{-\mathsf{x}-q_{\text{\tiny NS}}}{q_{\text{\tiny NS}}\mathsf{x}+1},~q_{\text{\tiny NS}}\neq 1$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho(\mathsf{x}))$
2 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\mp}$ $\mathsf{x}^{-1}$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}^{g_{\text{\tiny B}}}_{\tau}(\rho(\mathsf{x}))$
3 $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$ $\mathsf{x}^{-1},~~q_{\text{\tiny R}}=0$ ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho(\mathsf{x}))$
4 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$
Table 1: The inequivalent Möbius involutions $\rho(\mathsf{x})$.
Specifically, one finds that there is no solution $\rho_{4}(\mathsf{x})$ fixing the poles and permuting the zeros. Yet both poles and zeros are fixed by $\rho_{3}(\mathsf{x})=\mathsf{x}^{-1}$ when $q_{\text{\tiny R}}=0$. However, in this case the Lax connection does not depend on $\mathsf{x}$ and for trivial $K_{\sigma_{\star}}$ one has $T_{\text{\tiny b}}(\mathsf{x})=\mathbb{1}$ and trivial boundary conditions (4) that can not provide solutions of the sigma-model boundary equations of motion (LABEL:eq:sm-bcs).
The interesting cases arise when poles are permuted: Branch 1. Firstly, we have
$\rho_{1}(\mathsf{x})=-\frac{\mathsf{x}+q_{\text{\tiny NS}}}{q_{\text{\tiny NS}}\mathsf{x}+1},\qquad q_{\text{\tiny NS}}\neq 1,$ (10)
which fixes $\mathsf{x}_{0}$. For $q_{\text{\tiny NS}}=0$, this reduces to the standard reflection $\rho_{1}(\mathsf{x})=-\mathsf{x}$ obtained from parity-invariant gluing in pure RR cosets [7, 30, *Linardopoulos:2022wol, *Demjaha:2025axy]. Mapping residues of the conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho_{1}(\mathsf{x}))$ in general now gives
$\mathcal{J}_{+}^{\text{\tiny B}}\overset{\text{\tiny b}}{=}\frac{1-q_{\text{\tiny NS}}}{1+q_{\text{\tiny NS}}}\mathcal{J}^{\text{\tiny B}}_{-}~~\Rightarrow~~\partial_{\sigma}X^{\mu}\overset{\text{\tiny b}}{=}-q_{\text{\tiny NS}}\partial_{\tau}X^{\mu},$ (11)
[…]
$\partial_{\sigma}\tilde{X}^{\mu_{+}}\overset{\text{\tiny b}}{=}0,\qquad\partial_{\tau}\tilde{X}^{\mu_{-}}\overset{\text{\tiny b}}{=}0,$ (12)
where $\mu_{\pm}$ has cardinality $m_{\pm}$. For the automorphism given in (39), one finds e.g. $m_{+}=4$ and then this corresponds to a D-brane with $p=3$ and ${\cal F}=0$. Nontrivial worldvolume fluxes are typically introduced by dynamical and $\mathsf{x}$-dependent $K_{\sigma_{\star}}$-matrices, as happens for several interesting D-branes in $AdS_{5}\times S^{5}$ and $AdS_{4}\times\mathbb{CP}^{3}$ [30, *Linardopoulos:2022wol, *Demjaha:2025axy, 33]. In the future, it would be interesting to compare with the $\mathsf{x}$-dependent quantum $K_{\sigma_{\star}}$-matrices of [34, *Bielli:2024xuv, *Bielli:2025abu] known for pure RR flux. For nontrivial mixed-flux, an interesting question is also whether compatibility between (10) and (LABEL:eq:sm-bcs) requires dynamical reflection data $K_{\sigma_{\star}}$.
Branch 2. Both poles and zeros are permuted by
$\rho_{2}(\mathsf{x})=-\mathsf{x}^{-1},$ (13)
with no restriction on $q_{\text{\tiny NS}},q_{\text{\tiny R}}$. Including automorphisms $W$ and mapping residues, the minimal conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}W\left[{\cal L}^{g_{\text{\tiny B}}}_{\tau}(\rho_{2}(\mathsf{x}))\right]$ now simplify for $q_{\text{\tiny R}}\neq 0$ to
$\mathcal{J}^{\text{\tiny B}}_{\pm}\overset{\text{\tiny b}}{=}-W\left[\mathcal{R}^{\text{\tiny B}}_{\mp}\right]=-(W\mathrm{Ad}_{g_{\text{\tiny B}}})\left[\mathcal{J}^{\text{\tiny B}}_{\mp}\right],$ (14)
implying the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$. Interestingly, for $q_{\text{\tiny R}}=0$, where ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ become $\mathsf{x}$-independent, the boundary conditions enhance to $\mathcal{J}^{\text{\tiny B}}_{+}=-W\left[\mathcal{R}_{-}^{\text{\tiny B}}\right]$ without further constraints. These are precisely the conformal boundary conditions of WZW models which correspond to D-branes embedded as twisted conjugacy classes in $SU(1,1)\times SU(2)$. For $W=\mathbb{1}$ this yields unstable $dS_{2}\times S^{2}$ branes, which we will exclude in the following, while including the nontrivial outer automorphism (40) with $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ yields stable $AdS_{2}\times S^{2}$ branes [25]. In terms of the fields $X^{\mu}=\{\psi,\omega,\tau;\alpha,\beta,\gamma\}$ given in (36), the latter boundary conditions read
$\partial_{\tau}\psi$ $\overset{\text{\tiny b}}{=}0,\qquad$ $\partial_{\sigma}\omega$ $\overset{\text{\tiny b}}{=}\tanh\psi\cosh\omega\partial_{\tau}t,$ (15)
$\partial_{\sigma}t$ $\overset{\text{\tiny b}}{=}\tanh\psi\operatorname{sech}\omega\partial_{\tau}\omega,$
$\partial_{\tau}\alpha$ $\overset{\text{\tiny b}}{=}0,\qquad$ $\partial_{\sigma}\beta$ $\overset{\text{\tiny b}}{=}-\cot\alpha\sin\beta\partial_{\tau}\gamma,$
$\partial_{\sigma}\gamma$ $\overset{\text{\tiny b}}{=}\cot\alpha\csc\beta\partial_{\tau}\beta.$
The conjugacy classes are thus fixed at $\{\psi;\alpha\}$ and extend along $X^{m}=\{t,\omega;\beta,\gamma\}$. With (LABEL:eq:sm-bcs) and (37) the worldvolume flux can then be extracted from (15) as
${\cal F}_{mn}=\tfrac{\sinh 2\psi\cosh\omega}{2}\mathrm{d}\omega\wedge\mathrm{d}t+\tfrac{\sin 2\alpha\sin\beta}{2}\mathrm{d}\gamma\wedge\mathrm{d}\beta,$ (16)
For mixed-flux, however, the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$ fixes the D-brane conjugacy-classes on specific loci, i.e. at
$\psi=0,\qquad\alpha=\{0,\tfrac{\pi}{2},\pi\},$ (17)
on which the worldvolume flux becomes trivial, ${\cal F}=0$. We next show that this restriction can be lifted by allowing nontrivial reflection data (i.e. beyond $K_{\sigma_{\star}}=w_{\text{\tiny o}}$).
Before doing so, we remark that, interestingly, $\rho_{1}(\mathsf{x})$ at $q_{\text{\tiny NS}}=0$ can equally be interpreted as demanding the $\sigma\rightarrow 2\pi-\sigma$ parity-invariance of ${\cal L}(\mathsf{x})$ after reflection, while $\rho_{2}(\mathsf{x})$ combines requiring invariance under both parity and group element inversion $g_{\text{\tiny B}}\rightarrow g_{\text{\tiny B}}^{-1}$. It would be interesting to relate this to crossing properties of [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda].
Extension to all twisted conjugacy classes.
To restore the full family of stable $AdS_{2}\times S^{2}$ twisted conjugacy classes with two-form (16) at generic $(\psi,\alpha)$, we allow for appropriate matrices $K_{\sigma_{\star}}(\mathsf{x})$ while keeping the same reflection $\rho_{2}(\mathsf{x})=-\mathsf{x}^{-1}$. For this purpose we reverse the logic and use the WZW boundary conditions (15) as input to (4) to determine the required $K_{\sigma_{\star}}(\mathsf{x})$.
We find that it is sufficient to consider the ansatz
$K_{\sigma_{\star}}(\tau;\mathsf{x})=\left(\mathbb{1}+\sum_{k=1}^{6}u_{k}(X^{\mu};\mathsf{x})T_{k}\right)w_{\text{\tiny o}}\ ,$ (18)
with $T_{k}$ the $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ generators given in (27) and $u_{k}$ depending on $\tau$ only through the target-space fields $X^{\mu}=\{\psi,\omega,t;\alpha,\beta,\gamma\}$. Projecting the resulting system (4) onto the generators yields six coupled first-order PDEs in the worldsheet variables $(\tau,\sigma)$. Using (15), we rewrite them in terms of $\partial_{\sigma}\{\psi,\alpha\}$ and $\partial_{\tau}\{t,\omega,\beta,\gamma\}$ only. Requiring that $u_{k}$ does not depend on $\partial_{\sigma}\{\psi,\alpha\}$ (i.e. $u_{k}$ depends on the fields but not their derivatives), gives a set of algebraic conditions solved by
$\tfrac{u_{2}(\tau;\mathsf{x})}{u_{1}(\tau;\mathsf{x})}=\cot t,\quad$ $\tfrac{u_{3}(\tau;\mathsf{x})}{u_{1}(\tau;\mathsf{x})}=\tanh\omega\csc t,$ (19)
$\tfrac{u_{5}(\tau;\mathsf{x})}{u_{4}(\tau;\mathsf{x})}=-\cot\gamma,\qquad$ $\tfrac{u_{6}(\tau;\mathsf{x})}{u_{4}(\tau;\mathsf{x})}=-\cot\beta\csc\gamma.$
While this reduces the unfixed functions to $u_{1}$ and $u_{4}$ only, we remarkably find that the resulting six PDEs can still be solved consistently. We simply obtain
$u_{1}=-\sqrt{2}\frac{\sinh 2\psi\cosh\omega\sin t}{z(\mathsf{x})+\cosh 2\psi},~~u_{4}=\frac{\sqrt{2}}{i}\frac{\sin 2\alpha\sin\beta\sin\gamma}{z(\mathsf{x})-\cos 2\alpha},$ (20)
with the rational function $z(\mathsf{x})$ defined as
$z(\mathsf{x})=\frac{(\mathsf{x}-1)(q_{\text{\tiny NS}}(\mathsf{x}+1)+q_{\text{\tiny R}}(\mathsf{x}-1))}{(\mathsf{x}+1)((1-q_{\text{\tiny NS}})(\mathsf{x}-1)+q_{\text{\tiny R}}(\mathsf{x}+1))}.$ (21)
We obtain $u_{k}=0$ and hence $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ on the loci (17) and whenever the denominator of $z(\mathsf{x})$ vanishes, which occurs at $\mathsf{x}=\mathsf{x}_{0}$ and at $q_{\text{\tiny R}}=0$ (the WZW point). This is therefore fully consistent with the previous subsection. Away from these special points, we find that branes wrapping twisted conjugacy classes, as in boundary WZW models, remain compatible with integrability, with the mixed-flux deformation carried entirely by the nontrivial reflection data $K_{\sigma_{\star}}(\mathsf{x})$ rather than the brane geometry.
We will report on the extension to the fermionic sector, including the analysis of residual symmetries from $T_{\text{\tiny b}}(\mathsf{x})$ and the open-string spectrum, elsewhere.
Acknowledgements.
We thank D. Bielli, C. Kristjansen, S. Lacroix, G. Linardopoulos, B. Nairz, D. Polvara, A. Retore, and K. Zarembo for useful discussions, and A. Retore, K. Zarembo and especially S. Lacroix for comments on the draft. This work is partly based on the Master thesis of JCG prepared at ETH Zürich and we thank N. Beisert for acting as co-advisor. JCG’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”. SD is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and the NCCR SwissMAP. SD also thanks the organisers and participants of the workshop Higher-$d$ integrability (Favignana, 2025) where part of this work was presented.
Appendix A Sigma-model boundary conditions
We formulate integrable boundary conditions through a relation on the Lax connection (4), which guarantees conservation of charges generated by $T_{\text{\tiny b}}(\mathsf{x})$. At the level of the sigma-model, this must translate into boundary conditions on the fields $X^{\mu}$, $\mu=0,\ldots,D-1$, that are compatible with the variational principle.
The bosonic action of a standard two-derivative sigma-model on $\Sigma=\mathbb{R}_{\tau}\times[0,\pi]_{\sigma}$ can be written as  [^9] Here we use the flat sigma-model metric $\eta_{\alpha\beta}=\mathrm{diag}(1,-1)$, with $\sigma^{\alpha}=(\tau,\sigma)$, and $\epsilon_{\tau\sigma}=1$. In general the target space fields (metric and B-field) are curved, $G_{\mu\nu}:=G_{\mu\nu}(X)$ and $B_{\mu\nu}:=B_{\mu\nu}(X)$ and $T$ is the sigma-model coupling (i.e. string tension). As an action of open strings (22) is written in conformal gauge.
$S={}$ $\frac{T}{2}\int_{\Sigma}\mathrm{d}^{2}\sigma\;\partial_{\alpha}X^{\mu}\big(\eta^{\alpha\beta}G_{\mu\nu}+\epsilon^{\alpha\beta}B_{\mu\nu}\big)\partial_{\beta}X^{\nu}$ (22)
$+\int_{\partial\Sigma}\mathrm{d}\tau\;A_{m}\,\partial_{\tau}X^{m},$
where $m,n=0,\cdots,p\leq D-1$ are the directions along which $X^{\mu}|_{\partial\Sigma}$ is allowed to fluctuate, i.e. (for open strings) those along a Dpp-brane with a $U(1)$ gauge field $A_{m}:=A_{m}(X)$ and field strength $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$. The boundary equations of motion are then solved by
Dirichlet: $\partial_{\tau}X^{\hat{m}}\overset{\text{\tiny b}}{=}0,$ (23)
gen. Neumann: $G_{mn}\partial_{\sigma}X^{n}\overset{\text{\tiny b}}{=}{\cal F}_{mn}\partial_{\tau}X^{n},$
with $\hat{m}=p+1,\cdots,D-1$ and ${\cal F}_{mn}(X)=B_{mn}(X)+T^{-1}F_{mn}(X)$ the gauge-invariant worldvolume flux. Here we assumed adapted coordinates in which the target-space fields split orthogonally in the Neumann ($m$) and Dirichlet directions ($\hat{m}$). Note that the conditions (LABEL:eq:sm-bcs) are rather restrictive, given the symmetry properties of each term in the generalised Neumann conditions.
The problem addressed in this Letter can be restated as determining those Dirichlet and Neumann conditions which are guaranteed to preserve classical integrability. For the sigma-models considered here, the Lax connection ${\cal L}(\mathsf{x})\in\mathbb{C}(\mathsf{x})\otimes\mathfrak{g}$ is a linear combination of (projections of) $\mathfrak{g}$-valued one-form currents $\cal J$} and $\star{\cal J}$ with $\mathsf{x}$-dependent coefficients. The currents themselves are linear in $\mathrm{d}X^{\mu}$ so that one can introduce a frame field $E$ as
$\mathcal{J}_{\alpha}=E_{\mu}{}^{A}(X)T_{A}\partial_{\alpha}X^{\mu},\quad(\star\mathcal{J})_{\alpha}=\mathcal{J}_{\beta}\epsilon^{\beta}{}_{\alpha}$, (24)
with $T_{A}$ the generators of $\mathfrak{g}$. Imposing (4) for chosen $(\rho,U,K_{\sigma_{\star}})$ therefore yields linear gluing relations between $\partial_{\sigma}X^{\mu}$ and $\partial_{\tau}X^{\mu}$ at $\partial\Sigma$ which must be matched to the expressions (LABEL:eq:sm-bcs) for consistency with the variational principle, and hence determine the dimension $p$ and the flux $\mathcal{F}_{mn}$ (equivalently $F_{mn}$ for given $B_{mn}$). Likewise, one may start from local D-brane data $(p,{\cal F})$, and solve (4) for $(\rho,U,K_{\sigma_{\star}})$ using the relations (LABEL:eq:sm-bcs).
Appendix B Mixed-flux $AdS_{3}\times S^{3}$ bulk sigma-model
In this appendix, we collect the relevant ingredients for the boundary integrability analysis of the mixed NSNS/RR flux string on $AdS_{3}\times S^{3}\times M_{4}$, with $M_{4}$ a four-dimensional compact manifold required to complete the type IIB supergravity background (see [12, *Seibold:2024qkh] for reviews). Since the integrable structure relevant for our purposes is encoded in the $AdS_{3}\times S^{3}$ sector, we will restrict attention to its Green-Schwarz formulation as a $\mathbb{Z}_{4}$-graded supercoset sigma-model [14] (see also [37, *Hoare:2014oua]), while keeping $M_{4}$ implicit throughout.
Supercoset.
Concretely, the worldsheet degrees of freedom in this sector are described by $\widehat{F}/F^{(0)}$ with
$\widehat{F}=G\times G,\qquad G=PSU(1,1|2),$ (25)
and $F^{(0)}$ the bosonic diagonal subgroup of $\widehat{F}$, i.e. $(h,h)\in F^{(0)}$ for $h\in SU(1,1)\times SU(2)$. The Lie superalgebra $\hat{\mathfrak{f}}=\mathfrak{g}\oplus\mathfrak{g}$, with $\mathfrak{g}=\mathfrak{psu}(1,1|2)$, of $\widehat{F}$ admits a $\mathbb{Z}_{4}$ grading
$\hat{\mathfrak{f}}=\mathfrak{f}^{(0)}\oplus\mathfrak{f}^{(1)}\oplus\mathfrak{f}^{(2)}\oplus\mathfrak{f}^{(3)},$ (26)
with $\mathfrak{f}^{(0)}$ the Lie algebra corresponding to $F^{(0)}$. Here $\mathfrak{g}=\mathfrak{psu}(1,1|2)$, whose bosonic subalgebra $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ is generated by
$\begin{gathered}T_{1,2}=\tfrac{1}{\sqrt{2}}\left(\sigma_{1,3}\oplus 0\right),\qquad T_{3}=\tfrac{i}{\sqrt{2}}\left(\sigma_{2}\oplus 0\right),\\ T_{4,5,6}=\tfrac{1}{\sqrt{2}}\left(0\oplus\sigma_{1,2,3}\right),\end{gathered}$ (27)
with $\sigma_{1,2,3}$ the Pauli matrices. For a complete superalgebra realisation of $\mathfrak{g}$} and $\hat{\mathfrak{f}}$, see e.g. App. C of [39].
[…]
$\mathfrak{J}^{(0)}$ $=\begin{pmatrix}\mathcal{J}^{(0)}&\\ &\mathcal{J}^{(0)}\end{pmatrix}$$\quad$ $\mathfrak{J}^{(1)}=\begin{pmatrix}\mathcal{J}^{(1)}&\\ &-i\mathcal{J}^{(1)}\end{pmatrix}$, (28)
$\mathfrak{J}^{(2)}$ $=\begin{pmatrix}\mathcal{J}^{(2)}&\\ &-\mathcal{J}^{(2)}\end{pmatrix}$$\quad$ $\mathfrak{J}^{(3)}=\begin{pmatrix}\mathcal{J}^{(3)}&\\ &i\mathcal{J}^{(3)}\end{pmatrix}$,
[…]
$\operatorname{STr}(\mathfrak{J})=\mathrm{Str}(\mathcal{J})+\mathrm{Str}(\widetilde{\mathcal{J}}),\quad\widetilde{\operatorname{STr}}(\mathfrak{J})=\operatorname{STr}(\mathfrak{W}\mathfrak{J}),$ (29)
with $\mathrm{Str}$ the supertrace of $\mathfrak{g}$ and $\mathfrak{W}=\mathrm{diag}(\mathbb{1},-\mathbb{1})$.
Action and integrability.
The bulk sigma-model action with mixed-flux contributions is [14, 37, *Hoare:2014oua]
$S_{\text{\tiny cl.}}={}$ $\frac{1}{2}\int_{\Sigma}\operatorname{STr}\left(\mathfrak{J}^{(2)}\wedge\star\mathfrak{J}^{(2)}+q_{\text{\tiny R}}\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(3)}\right)$ (30)
$+q_{\text{\tiny NS}}\int_{\cal B}\widetilde{\operatorname{STr}}\left(\frac{2}{3}\mathfrak{J}^{(2)}\wedge\mathfrak{J}^{(2)}\wedge\mathfrak{J}^{(2)}\right$.
$+\left.\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(3)}\wedge\mathfrak{J}^{(2)}+\mathfrak{J}^{(3)}\wedge\mathfrak{J}^{(1)}\wedge\mathfrak{J}^{(2)}\right),$
[…]
Classical integrability, $\kappa$-symmetry and conformal invariance to one-loop in $\alpha^{\prime}$ require the following relation between the flux parameters [14]
$q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1.$ (31)
The flat Lax connection then takes the form
${\cal L}(\mathsf{x})={}$ ${\cal J}^{(0)}+q_{\text{\tiny R}}\frac{\mathsf{x}^{2}+1}{\mathsf{x}^{2}-1}\mathcal{J}^{(2)}+\left(q_{\text{\tiny NS}}-\frac{2q_{\text{\tiny R}}\mathsf{x}}{\mathsf{x}^{2}-1}\right)\star\mathcal{J}^{(2)}$
$+\left(\mathsf{x}+\frac{q_{\text{\tiny R}}}{1-q_{\text{\tiny NS}}}\right)\sqrt{\frac{q_{\text{\tiny R}}(1-q_{\text{\tiny NS}})}{\mathsf{x}^{2}-1}}\mathcal{J}^{(1)}$ (32)
$+\left(\mathsf{x}-\frac{q_{\text{\tiny R}}}{1+q_{\text{\tiny NS}}}\right)\sqrt{\frac{q_{\text{\tiny R}}(1+q_{\text{\tiny NS}})}{\mathsf{x}^{2}-1}}\mathcal{J}^{(3)}.$
At the pure RR point $q_{\text{\tiny R}}=1$, $q_{\text{\tiny NS}}=0$ this Lax corresponds to [40], while at the pure NSNS/WZW point it degenerates and does not evidently capture the full dynamics. In that case one can instead work with holomorphic current algebras rather than Lax connections.
Parametrisation and background.
The sigma-model (22) is invariant under local, right-acting, $F^{(0)}$ gauge transformations. We can therefore take
$g=g_{\text{\tiny F}}g_{\text{\tiny B}},\qquad\tilde{g}=\tilde{g}_{\text{\tiny F}}\tilde{g}_{\text{\tiny B}}=\tilde{g}_{\text{\tiny F}},$ (33)
[…]
$\mathcal{J}=\mathcal{J}^{\text{\tiny B}}+g_{\text{\tiny B}}^{-1}\mathcal{J}^{\text{\tiny F}}g_{\text{\tiny B}},\qquad\widetilde{\mathcal{J}}=\widetilde{\mathcal{J}}^{\text{\tiny F}},$ (34)
with $\mathcal{J}^{\text{\tiny B}}=g_{\text{\tiny B}}^{-1}\mathrm{d}g_{\text{\tiny B}}\in\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$, $\mathcal{J}^{\text{\tiny F}}=g_{\text{\tiny F}}^{-1}\mathrm{d}g_{\text{\tiny F}}\in\mathfrak{g}$ and $\widetilde{\mathcal{J}}^{\text{\tiny F}}=\tilde{g}_{\text{\tiny F}}^{-1}\mathrm{d}\tilde{g}_{\text{\tiny F}}\in\mathfrak{g}$. In the main text, we furthermore focus on the bosonic sector of $\mathfrak{psu}(1,1|2)$ and thus take $g_{\text{\tiny F}},\tilde{g}_{\text{\tiny F}}=\mathbb{1}$. Altogether, this implies the simplifications
$\left.\mathcal{J}^{(0)}\right|_{g_{\text{\tiny B}}}=\left.\mathcal{J}^{(2)}\right|_{g_{\text{\tiny B}}}=\frac{1}{2}\mathcal{J}^{\text{\tiny B}},\qquad\left.\mathcal{J}^{(1,3)}\right|_{g_{\text{\tiny B}}}=0.$ (35)
We will then parametrise the bosonic element $g_{B}=\mathrm{diag}(g_{SU(1,1)}~|~g_{SU(2)})$ as
$g_{SU(1,1)}$ $=\sinh\psi~\sigma_{3}+\cosh\psi~e^{\tfrac{it}{2}\sigma_{2}}e^{\omega\sigma_{1}}e^{\tfrac{it}{2}\sigma_{2}},$ (36)
$g_{SU(2)}$ $=e^{-\tfrac{i\gamma}{2}\sigma_{3}}e^{\tfrac{i\beta}{2}\sigma_{1}}e^{i\alpha\sigma_{3}}e^{-\tfrac{i\beta}{2}\sigma_{1}}e^{\tfrac{i\gamma}{2}\sigma_{3}},$
where ${\psi,\omega,t}\in\mathbb{R}^{3}$ are global coordinates on $AdS_{3}$, with $\{\omega,t\}$ parametrising $AdS_{2}$ fibred over $\psi$, while the coordinates $\beta\in[0,\pi)$ and $\gamma\in[0,2\pi)$ parametrise an $S^{2}$ fibered over the interval $\alpha\in[0,\pi)$, forming the three-sphere $S^{3}$. The metric and NSNS three-form $H=\mathrm{d}B$ can then be expressed as
$\mathrm{d}s^{2}={}$ $\mathrm{d}\psi^{2}+\cosh^{2}\psi\left(-\cosh^{2}\omega\mathrm{d}t^{2}+\mathrm{d}\omega^{2}\right)$ (37)
$+\mathrm{d}\alpha^{2}+\sin^{2}\alpha\left(\mathrm{d}\beta^{2}+\sin^{2}\beta\mathrm{d}\gamma^{2}\right),$
$H={}$ $2q_{\text{\tiny NS}}\cosh^{2}\psi\cosh\omega\mathrm{d}\psi\wedge\mathrm{d}t\wedge\mathrm{d}\omega$
$+2q_{\text{\tiny NS}}\sin^{2}\alpha\sin\beta\mathrm{d}\alpha\wedge\mathrm{d}\beta\wedge\mathrm{d}\gamma.$
When including the fermionic sector, the RR 3-form flux is $F_{3}=e^{-\Phi_{0}}q_{\text{\tiny R}}(H/q_{\text{\tiny NS}})$ for constant dilaton $\Phi(X)=\Phi_{0}$.
Outer automorphisms of $\bm{SU(1,1)\times SU(2)}$.
Automorphisms $W$ of $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ which preserve the ad-invariant metric form the group $SO(2,1)\times SO(3)$. Explicit matrix realisations can be defined as $W(y)=wyw^{-1}=y^{A}W_{A}{}^{B}T_{B}$ for $y=y^{A}T_{A}\in\mathfrak{g}$. The elements relevant to the boundary integrability discussion of the main text are those $W$ which are involutive, $W^{2}=\mathbb{1}$. Up to conjugation (and combinations), the ones that do not mix $\mathfrak{su}(1,1)$ and $\mathfrak{su}(2)$ and that (i) are inner are
$W_{\text{\tiny inn}}=\left\{\begin{array}[]{ll}\mathbb{1}_{3}\oplus\mathbb{1}_{3}\\ \mathrm{diag}(-1,-1,1)\oplus\mathbb{1}_{3}\\ \mathbb{1}_{3}\oplus\mathrm{diag}(-1,-1,1)\end{array}\right.~\in SO^{+}(2,1)\times SO(3),$ (38)
or (ii) are outer are
$W_{\text{\tiny out.}}=\mathrm{diag}(-1,1,-1)\oplus\mathbb{1}_{3}\in SO^{-}(2,1)\times SO(3).$ (39)
The latter is realised as $W_{\text{\tiny out.}}(X)=w_{\text{\tiny o}}Xw_{\text{\tiny o}}^{-1}$ with
$w_{\text{\tiny o}}=\mathrm{diag}(1,-1)\oplus\mathbb{1}_{2}\not\in SU(1,1)\times SU(2).$ (40)